Classify each equation as linear or nonlinear dy/dx = y^3 - 9 linear y = 3y - 6 nonlinear > 3y y = 6-4 linear > 4x^2 y -
Answer #1

We know that general equation of linear differerential equation is an(x) any + Cu(a) dy + Q(x) yg = E(X) any (a) do dan y dany=34-6 12y and get and y = on da? Put y= dy da 3dy - 6 day dx2 da this equation is linear 422 da? 3 dy =-6. da 34 y=64-4